OpenStudy (mony01):

Someone know how to isolate c from -12/(c-3)^3

OpenStudy (yttrium):

We can't isolate the c here if there is no equation. Or we can do that when we substitute values for c.

OpenStudy (mony01):

here is the question @Yttrium Let f(x) = (x − 3)^−2. Find all values of c in (1, 7) such that f(7) − f(1) = f '(c)(7 − 1).

OpenStudy (yttrium):

any help @hartnn

hartnn (hartnn):

which topic does this belong ? mean value theorem ?

OpenStudy (mony01):

yeah

hartnn (hartnn):

could you find f'(x) first ?

hartnn (hartnn):

from f'(x) you could get f'(c), by just plugging in x=c

OpenStudy (mony01):

isn't f'(x)=-2(x-3)^-3

hartnn (hartnn):

correct, so f'(c) = -2 (c-3)^-3 now find f(7)and f(1) and plug those in f(7) − f(1) = f '(c)(7 − 1).

OpenStudy (mony01):

I did and i got -12/(c-3)^3

hartnn (hartnn):

you should get an equation...

OpenStudy (mony01):

For f(7)=1/16 and f(1)=1/4

OpenStudy (mony01):

so is the c= -3/16

hartnn (hartnn):

no, f(7) and f(1) were incorrectly calculated.

OpenStudy (yttrium):

Oww. I thougt f'(c) --> Inverse function of x wrt c. :((

OpenStudy (mony01):

dont you have to calculate for c?

hartnn (hartnn):

to get f(7), just plug in x=7 in f(x) 'c' doesn't even come in picture to get f(7) and f(1)

OpenStudy (mony01):

oh okay but to calculate this part how do you do it f '(c)(7 − 1).

hartnn (hartnn):

what did you get for f(7) and f(1) ?

OpenStudy (mony01):

together -3/16

hartnn (hartnn):

what is together ? f(7) = (x-3)^-2 at x=7 = 1/16 f(1) = 1/4 you were correct about f(7) and f(1) so, f(7) − f(1) = f '(c)(7 − 1) gives 1/16 -1/4 = -2(c-3)^-3 *6 just try to isolate 'c' from here

hartnn (hartnn):

so, -3/16 = -12 (c-3)^-3 1/64 = 1/(c-3)^3 gives, 64 = (c-3)^3 take cube root on both sides, what u get ?

OpenStudy (mony01):

is it 7

hartnn (hartnn):

yep, c=7, thats correct :)

OpenStudy (mony01):

is there any more because when I put 7 I got it wrong

hartnn (hartnn):

yes, thats because c has the range (1,7) which means 1 and 7 are not included so, c cannot take the value of 7

OpenStudy (mony01):

so it does not exist

hartnn (hartnn):

yup, i would say there's no solution

OpenStudy (mony01):

which one would you say it is from the attach list?

hartnn (hartnn):

i would have gone with b as MVT applies for continuous functions

OpenStudy (mony01):

cool got them both right thank you so much hartnn!!!!!!

hartnn (hartnn):

nice ! welcome ^_^